OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.D(5, -2) m = 2/5

OpenStudy (jdoe0001):

http://www.teachhandson.com/calculationb.gif use the slope given, use the point given \(\large (x_2, y_2)\) plug them in the point-slope form, and solve for "y" to get the so-called "standard form"

terenzreignz (terenzreignz):

Okay, we have a few numbers here...

terenzreignz (terenzreignz):

The slope m and the two numbers given by the point (a,b) Right? :D

OpenStudy (anonymous):

I'm really bad at math and any time my brother tries to explain it i get frustrated and storm off. i really need help as simple as possible.

terenzreignz (terenzreignz):

Yes. It will be simple. We have three numbers, the slope m and the point (a,b) [a and b being the numbers I describe] Catch me so far?

OpenStudy (anonymous):

i think so.

terenzreignz (terenzreignz):

Okay, so to get the line passing through a point (a,b) and having a slope m it has the equation... \[\Large y = \color{green}m(x-\color{red}a)+\color{blue}b\]

terenzreignz (terenzreignz):

So, just fill in the coloured values (m, a, and b) with their appropriate numbers... For instance, the slope m is 2/5 So replace m with 2/5 \[\Large y = \color{green}{\frac25}(x-\color{red}a)+\color{blue}b\]

terenzreignz (terenzreignz):

What are your values for a and b?

OpenStudy (anonymous):

5 and -2?

terenzreignz (terenzreignz):

That's right :D \[\large \color{red}a = \color{red}5\]\[\large \color{blue}b = \color{blue}{-2}\] So replace a and b accordingly \[\Large y = \color{green}{\frac25}(x-\color{red}5)+\color{blue}{(-2)}\]

terenzreignz (terenzreignz):

And now just simplify :)

OpenStudy (anonymous):

wait so whats 2/5(x-5)?

terenzreignz (terenzreignz):

You distribute the 2/5 over the (x - 5) You know, like this... \[\Large p(q+r) = pq+pr\]

OpenStudy (jhannybean):

p(q-r) *

terenzreignz (terenzreignz):

tch same concept \[\Large p(q-r) = pq-pr\] :P

terenzreignz (terenzreignz):

Btw don't you think this belongs in Algebra and not Geometry? @Jhannybean ? :D

OpenStudy (anonymous):

it probably should but this is a part of my geometry homework and im way behind :/

OpenStudy (jhannybean):

Guess so, but sometimes people get confused which section to post their questions in, or because the general "math" section has more people responding to posts.

terenzreignz (terenzreignz):

Okay, no worries guys :D